Write a quadratic equation in the variable x having the given numbers as solutions.Type the equation in standard form.
Solutions -radical 5, 7 radical 5

Answer by ewatrrr(24785)

y = (x + √5)(x - 7√5)
(x + √5)(x - 7√5) = 0
x^2 -(6√5)x - 35 = 0
F First terms
O Outside terms
I Inside terms
L Last terms √5(-7√5) = -7•5